OpenStudy (anonymous):

a man weight 1645n,what is his mass

OpenStudy (aravindg):

Weight=Mass x g=Mass x 9.8 Mass=weight/9.8

OpenStudy (gebooors):

Weight G = mg, in which g = 9,80665 m/s\[^{2}\] m = G/g = 1654 N / 9,80665 m/s2 = 168,66106 kg Also with exact accuracy mass is 168,7 kg. Notice that you have to use more accurate value for g as usual.

OpenStudy (anonymous):

M = w/g = 167.7kg

OpenStudy (chmvijay):

Do u mean "his mass" on earth or moon or Jupiter bcoz g varies with planet :) for the earth above mentioned is correct :)